### v1

@@ Abstract
The development of large-scale 3D scene reconstruction and novel view synthesis methods mostly rely on datasets comprising perspective images with narrow fields of view (FoV). While effective for small-scale scenes, these datasets require large image sets and extensive structure-from-motion (SfM) processing, limiting scalability. To address this, we introduce a fisheye image dataset tailored for scene reconstruction tasks. Using dual 200-degree fisheye lenses, our dataset provides full 360-degree coverage of 5 indoor and 5 outdoor scenes. Each scene has sparse SfM point clouds and precise LIDAR-derived dense point clouds that can be used as geometric ground-truth, enabling robust benchmarking under challenging conditions such as occlusions and reflections. While the baseline experiments focus on vanilla Gaussian Splatting and NeRF based Nerfacto methods, the dataset supports diverse approaches for scene reconstruction, novel view synthesis, and image-based rendering.
@@ 3.2 Data Collection
To obtain the geometry ground truth of each scene, we used the Faro Focus 3D LIDAR scanner fixed on a tripod, to capture high-resolution XYZRGB point clouds. The Faro scanner covers a 360°horizontal and 170°vertical FoV ($-60$°to 90°) with a data capture range of 0.6–200 meters depending on the indoor or outdoor capture modes. Similar to the camera setup, the Faro scanner was placed at fixed locations in the scene and multiple scans, each corresponding to a different location in the scene, are taken at 1/4 or 1/5 resolution and 4×4$\times$4 × quality. Each scan lasted around 11 minutes and depending on the scale of the scene, 5–25 scans at different fixed positions were necessary to fully cover each scene. Minor artifacts from moving objects were negligible relative to scene scale and were not visible in the dense point clouds created. The dense point cloud captured with the Faro Scanner for the scene labeled as Kitchen_In is given in Figure 1 as an example. Representative visuals from the dense point clouds of our scenes are provided in Supplementary Materials.
@@ 3.3 Formation of Sparse and Dense Point Clouds
The Structure-from-Motion (SfM) is a fundamental step in many 3D scene reconstruction methods, as it allows for the recovery of camera poses and the forming of sparse 3D structures from multiple overlapping images. In our dataset, we utilized COLMAP [27][29] version 3.9.1, an incremental SfM pipeline, to generate sparse point clouds of the scenes using the raw fisheye images and their calibrated parameters.
@@
The COLMAP SfM pipeline outputs the sparse point cloud representing the 3D structure of each scene in binary (.bin) format. The binary format is included in each scenes’ sparse model in our dataset, and these can be converted to the text format if necessary. Additionally, PLY-format point clouds are included in the dataset for easier visual verification, as well as for an easier alignment process with denser ground-truth point clouds, and use by other researchers. The COLMAP format is also supported by Nerfstudio [32], a common 3D scene reconstruction framework that can enable real-time rendering in navigable environments for a number of NeRF and 3DGS-based scene reconstruction methods.
@@ 3.5 Image Rectification
The rectification process was implemented for convenience and allowed the dataset to be easily integrated into the novel 3D scene reconstruction pipelines. However, this approach is not optimal, as the rectified images may introduce slight distortions or inaccuracies, and lose scene information from the heavily distorted parts of the fisheye images. Despite this, the processed dataset provides a practical solution for our experiments and works sufficiently well, based on the quantitative results we obtain.
@@ 4 Experiments
In this section, we present two experiments that demonstrate the capabilities of our dataset, highlighting the usage of both the sparse COLMAP data and the dense Faro scanner data. For the experiments conducted here, we applied a standard 90%-10% train-test split to the images for each scene. The visual results and evaluation metrics presented correspond to the rendered test images, which were randomly selected from the full set of images for each scene.
@@
In the first experiment, we establish a baseline for rendering quality and performance using the Gaussian splatting (3DGS) [14] and Nerfstudio (v1.1.4)’s NeRF-based scene reconstruction method Nerfacto [32]. For both of these models we use the sparse point clouds generated by the SfM software COLMAP, which are provided with the dataset. These point clouds originate from fisheye images, which we rectify (undistort) using camera calibration parameters to be compatible with the Gaussian splatting implementation’s rasterization requirements for pinhole camera models [14], and with the Nerfacto models’ supported camera types.
@@ 5 Conclusion
We introduced a high-quality dataset to address limitations in existing resources for large-scale 3D scene reconstruction, novel view synthesis, and image-based rendering. Captured using an Insta360 camera with dual 200-degree fisheye lenses, the dataset provides comprehensive 360-degree coverage while compensating for heavy lens distortion through careful calibration and maintaining high scene detail. Complemented by high-precision ground truth point clouds from a Faro Focus 3D LiDAR scanner, it enables robust geometric evaluation and alignment benchmarking. The dataset presents unique challenges, such as diverse weather conditions, occlusions, and reflective surfaces, making it well suited to studying algorithm robustness under real-world complexities. By relying on still images, it avoids motion blur, maintains high detail and provides a solid basis for advancing 3D reconstruction and novel view synthesis in complex environments.
@@ 6 Acknowledgements
This work was supported by IWM Doctoral Pilot, Ministry of Education and Culture, Finland, and the CIVIT Research Infrastructure, Finland.

### v2

@@ Abstract
The development of large-scale 3D scene reconstruction and novel view synthesis methods mostly rely on datasets comprising perspective images with narrow fields of view (FoV). While effective for small-scale scenes, these datasets require large image sets and extensive structure-from-motion (SfM) processing, limiting scalability. To address this, we introduce a fisheye image dataset tailored for scene reconstruction tasks. Using dual 200-degree fisheye lenses, our dataset provides full 360-degree coverage of 5 indoor and 5 outdoor scenes. Each scene has sparse SfM point clouds and precise LIDAR-derived dense point clouds that can be used as geometric ground-truth, enabling robust benchmarking under challenging conditions such as occlusions and reflections. While the baseline experiments focus on vanilla Gaussian Splatting and NeRF based Nerfacto methods, the dataset supports diverse approaches for scene reconstruction, novel view synthesis, and image-based rendering. The dataset is available here.
@@ 3.2 Data Collection
To obtain the geometry ground truth of each scene, we used the Faro Focus 3D LIDAR scanner fixed on a tripod, to capture high-resolution XYZRGB point clouds. The Faro scanner covers a 360°horizontal and 170°vertical FoV ($-60$°to 90°) with a data capture range of 0.6–200 meters depending on the indoor or outdoor capture modes. Similar to the camera setup, the Faro scanner was placed at fixed locations in the scene and multiple scans, each corresponding to a different location in the scene, are taken at 1/4 or 1/5 resolution and 4×4$\times$4 × quality. Each scan lasted around 11 minutes and depending on the scale of the scene, 5–25 scans at different fixed positions were necessary to fully cover each scene. Minor artifacts from moving objects were negligible relative to scene scale and were not visible in the dense point clouds created. An example dense point cloud captured with the Faro Scanner for the Kitchen_In scene is given in Figure 1.
@@ 3.3 Formation of Sparse and Dense Point Clouds
The Structure-from-Motion (SfM) is a fundamental step in many 3D scene reconstruction methods, as it allows for the recovery of camera poses and the forming of sparse 3D structures from multiple overlapping images. In our dataset, we utilized COLMAP [27][29] version 3.9.1, an incremental SfM pipeline, to generate sparse point clouds of the scenes using the raw fisheye images.
@@
The COLMAP SfM pipeline outputs the sparse point cloud representing the 3D structure of each scene in binary (.bin) format. The binary format is included in each scenes’ sparse model in our dataset, and these can be converted to the text format if necessary. The COLMAP format is also supported by Nerfstudio [32], a common 3D scene reconstruction framework that can enable real-time rendering in navigable environments for a number of NeRF and 3DGS-based scene reconstruction methods.
@@ 3.5 Image Rectification
The rectification process was implemented for convenience and allowed the dataset to be easily integrated into the novel 3D scene reconstruction pipelines. However, this approach is not optimal, as the rectified images may lose scene information from the heavily distorted parts of the fisheye images. Despite this, the processed dataset provides a practical solution for our experiments and works sufficiently well, based on the quantitative results.
@@ 4 Experiments
In this section, we present two experiments that demonstrate the capabilities of our dataset, highlighting the usage of both the sparse COLMAP data and the dense Faro scanner data. For our experiments, we applied a standard 90%-10% train-test split to the images for each scene. The visual results and evaluation metrics presented correspond to the rendered test images, which were randomly selected from the complete set of images for each scene.
@@
In the first experiment, we establish a baseline for rendering quality and performance using the Gaussian splatting (3DGS) [14] and Nerfstudio (v1.1.4)’s NeRF-based scene reconstruction method Nerfacto [32]. For both of these models we use the sparse point clouds generated by the SfM software COLMAP. These point clouds originate from fisheye images, which we rectify (undistort) using camera calibration parameters to be compatible with the Gaussian splatting implementation’s rasterization requirements for pinhole camera models [14], and with the Nerfacto models’ supported camera types.
@@ 5 Conclusion
We introduced a high-quality dataset to address limitations in existing resources for large-scale 3D scene reconstruction, novel view synthesis, and image-based rendering. Captured using an Insta360 camera with dual 200-degree fisheye lenses, the dataset provides comprehensive 360-degree coverage while compensating for heavy lens distortion through calibration and maintaining high scene detail. Complemented by dense ground truth point clouds from a Faro Focus 3D LiDAR scanner, it enables robust geometric evaluation and alignment benchmarking. The dataset presents unique challenges, which makes it well suited to 3D scene reconstruction under real-world complexities. By relying on still images, it avoids motion blur, maintains high detail, and provides a solid basis for advancing 3D reconstruction and novel view synthesis in complex environments.
@@ 6 Acknowledgements
We acknowledge the financial support of the Intelligent Work Machines Doctoral Education Pilot Program (IWM VN/3137/2024-OKM-4) and funding from the Research Council of Finland (grants 352788, 353138, 362407, 362408, 339730, 353139, 362409) and the Finnish Center for Artificial Intelligence. We also acknowledge the Centre for Immersive Visual Technologies for the equipment used.